The perimeter of a rectangular garden is 308 feet. If the length of the garden is 97 feet, what is its width?

Answer by Alan3354(69443)

The perimeter of a rectangular garden is 308 feet. If the length of the garden is 97 feet, what is its width?
----------
The perimeter of a rectangle is 2 x the length + 2 x the width.
1/2 the perimeter is L+W.
308/2 = 97 + W
154 = 97 + W
W = 57 feet